OpenStudy (anonymous):

A 50 kg person is standing on a bathroom scale in an elevator moving upward. a.) if the scale reads 56 kg, what is the magnitude of upward acceleration of the elevator (m/s^2)? b.) Now the elevator is traveling up while slowing down. If the magnitude of deceleration is the same as that in part (a), what is reading given by the scale?

OpenStudy (anonymous):

\[R = m(g+a)\] \[560 = 50 (10 +a)\] \[560 = 500 + 50a\] \[60 = 50a\] \[a = \frac{ 6 }{ 5 } = 1.2m/s^2\]

OpenStudy (anonymous):

b) travelling up, decelerating: N = m(g- 6/5) N = m*8.8 = 440 N (scale reading)

OpenStudy (anonymous):

Thanks Yahoo! =)

OpenStudy (anonymous):

@Algebraic - thanks for replying to the question. I think they're wanting the answer to be in kg, so would I just convert the 440 N to kg? Or do I use another formula to find it?

OpenStudy (anonymous):

@gifty91 Do u have the answer key with u

OpenStudy (anonymous):

@gifty91

OpenStudy (anonymous):

Not really.. I just get 5 attempts at a question, and after the 5th attempt, it'll only show me the correct answer.